Use this table or Appendix B to write balanced formation equations at standard conditions for each of the following compounds. (Apply fractional coefficients as needed. Include states-of-matter under STP conditions in your answer.)

(a) Ca3(PO4)2(s)

(b) HCN(g)

(c) HI(g)

(d) NaCl(s)
Answer #1

(a) 3 Ca(OH)2 (l) + 2 H3PO4 (l) = Ca(PO4)2 (s) + 6 H20 (l)

(b) 2 CH4 (g) + 2 NH3 (l) + O2 (g) = 2 HCN (g) + 6 H20 (g)

(c) 2 I2 (g) + N2H4 (l) = 4 HI (g) + N2 (g)

d) NaOH (l) + HCl (l) = NaCl (s) + H20 (l)